OpenStudy (anonymous):

when you multiply it all out, how many digits are in the answer to 2^350 x 3^3 x 5^351

OpenStudy (anonymous):

There are floor(log(2^350*3^3*5^351))+1 digits Which is floor(350*log(2) + 3*log(3) + 351*log(5))+1 digits = floor(352.13033..)+1 digits = 353 digits. The floor function rounds the number down to the nearest integer.

OpenStudy (anonymous):

log is log base 10 for this question.

OpenStudy (anonymous):

wow really?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

that is cool. i would have done something more prosaic and said \[2^{350}\times 5^{351}=2^{350}\times 5^{350}\times 2=10^{350}\times 2\]\]

OpenStudy (anonymous):

and \[2\times 3^3=54\]

OpenStudy (anonymous):

@mathboy that is quite nice. but i think i am getting a different answer or am i confused?

OpenStudy (anonymous):

The 3rd 2 in your equation should be a 5 i think, if that makes any difference..

OpenStudy (anonymous):

(and the 4th)

OpenStudy (anonymous):

ooh it should be \[5\times 3^3=135\] got it

OpenStudy (anonymous):

so it would be 135 digits?

OpenStudy (anonymous):

hmm am i still off by one?

OpenStudy (anonymous):

oh not not 135 digits

OpenStudy (anonymous):

it is \[135\times 10^{350}\] which has 353 digits

OpenStudy (anonymous):

someone else said 352 digits.

OpenStudy (anonymous):

Yep i think that's correct.

OpenStudy (anonymous):

10^1 has 2 digits though right, so 10^x has x+1 digits.

OpenStudy (anonymous):

that some else who said 352 digits on another copy of this question was me - and I was off by one digit . 353 digits is correct - wolfram confirms it: http://www.wolframalpha.com/input/?i=2^350+x+3^3+x+5^351